OpenStudy (anonymous):

isolate the a d=v*t+1/2*a*t^2

OpenStudy (anonymous):

subtract a vt on both sides to get \[d-vt = (1/2)at^2\] then multiply both sides by 2 to get \[2(d-vt) = at^2\] then divide both side by t^2 to get your answer